The measures of two angles of a triangle are in the ratio 2:3. If the third angle is 4 degrees larger than the larger of the other two angles, find the measure of an exterior angle at the third vertex.
Answer by lwsshak3(11628)

The measures of two angles of a triangle are in the ratio 2:3. If the third angle is 4 degrees larger than the larger of the other two angles, find the measure of an exterior angle at the third vertex.
***
let x=larger of 2 angles of a triangle
let y=smaller of 2 angles of a triangle
y/x=2/3
y=(2/3)x
x+4=exterior angle at the third vertex.
..
x+y+x+4=180
x+(2/3)x+x+4=180
(8/3)x=176
x=3(176)/8=66
x+4=70
exterior angle at the third vertex=70˚